OpenStudy (anonymous):

f(x)= (x + 3) * (x - 5)^2 verity zeros PLEASE HELP ASAP!

OpenStudy (anonymous):

PLEASE HELP!

OpenStudy (anonymous):

For zeros, f(x)=0

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i found this online, but is it correct? Part1: f(x) = (x + 3) * (x - 5)^2 Part2: x = -3 ====> x + 3 = 0 x = 5 ====> x - 5 = 0 ====> (x - 5)^2 = 0 f(x) = (x + 3) * (x - 5)^2

OpenStudy (anonymous):

f(x)=0 (x + 3) * (x - 5)^2=0 Either (x+3)=0 or (x-5)^2 =0

OpenStudy (anonymous):

Ya thats correct

OpenStudy (anonymous):

x=-3 and 5

OpenStudy (anonymous):

and thats the answer?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

Welcome buddy